Fermi National Laboratory
 

Accelerator Update for June 8, 2007 - June 11, 2007

 

Friday June 8

The day shift began with Operations monitoring the Tevatron (TeV) store 5480 (46.22E30) and the Recycler stash (220.8E10), with the Antiproton Source (Pbar) stacking (48.9E10), with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI and MiniBooNE taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams.

Operations terminated store 5480 at 12:03 PM and then put the TeV into shot setup.

At 2:29 PM, a Recycler expert asked Operations to swap out a control card in a MI CAMAC crate ($23). Unfortunately, this action tripped off some Recycler corrector powers supplies and caused the loss of the stash – 18mA. Fortunately, the antiprotons had already been loaded for the next store. Operators restored the crate and card.

Operations established store 5481 at 2:37 PM with an initial luminosity of 213.66E30.

A TeV expert began a Tevatron Electron Lens (TEL1) study at 2:44 PM.

Operations reported at 3:59 PM that the Recycler had stashed antiprotons three times during the day shift.

At 10:20 PM, a Linac Klystron RF station (KRF6) tripped off. It would not reset. Operators called in an expert. Operators managed to get the station on before the experts arrived and then the experts tuned the station. Linac resumed sending beam at 11:09 PM.

Operations reported at midnight that the Recycler had stashed three times during the evening shift.

Saturday June 9

The midnight shift began with Operations monitoring store 5481 (79.9E30) and the stash (143.3E10), with Pbar stacking (42.5E10), and with NuMI and MiniBooNE taking beam.

At 4:29 AM, NuMI tripped off due to a power supply trip (QF112). It reset.

At 6:18 AM, two MI kickers tripped off (I:KPS4A and KPS4B). Operators had them back on in about 10 minutes.

Operations terminated store 5481 at 8:02 AM.

Operations prepared D0 for an access at 8:12 AM. They came out of access at 9:36 AM.

Operations put the TeV through a dry squeeze at 10:10 AM.

Operations put the TeV into shot setup at 10:43 AM.

At 12:04 PM, Operations reported some problems extracting antiprotons from the Recycler. Operations called in an expert. The expert investigated and found that studiers on Friday had left the phase damping on after their bucket waveform work. Operators turned the damping off.

Operations established store 5483 at 12:56 PM with an initial luminosity of 164.77E30.

At 3:30 PM, a MI kicker (KPS1A) tripped off twice in one minute. Operations called in an expert.

At 6:33 PM, Operations reported that KPS1A had been tripping off every four or five minutes. The expert worked on the kicker. Pbar stacked in between the trips.

Operations reported at 8:03 PM that the Recycler had stashed two times during the last twelve hours. Accelerator status: store 5483 was at 72.48E30, the stack was at 27.3E10, and the stash was at 171.8E10.

At 9:15 PM, Operations reported that KPS1A had tripped off for the first of four times in 50 minutes. It reset.

Sunday June 10

The midnight shift began with Operations monitoring store 5483 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

Operations terminated store 5483 at 8:01 AM.

Operations reported at 8:02 AM that the Recycler had stashed five times during the last twelve hours. Accelerator status: the stack was at 37.3E10, and the stash was at 328.5E10.

Operations put the TeV into shot setup at 8:17 AM.

At 9:29 AM, Operations reported that they were having problems trying to close the TeV antiproton orbit. Operators contacted TeV experts.

Operations established store 5484 at 10:18 AM with an initial luminosity of 226.12E30.

At 10:21 AM, Operations reported that LRF3 had tripped off. It reset.

At 10:34 AM, Operations reported that LRF5 had tripped off. It reset.

At 1:38 PM, Operations reported that KPS1A had tripped off for the first of five times in 13 minutes. It reset.

Operations reported at 3:47 PM that to keep KPS1A up and running they had adjusted its reservoir and filament voltages.

At 4:05 PM, KPS1A tripped off two more times. Operators reset the kicker and then lowered the room temperature by three degrees.

Operations reported at 8:06 PM that the Recycler had stashed four times during the last twelve hours. Accelerator status: store 5484 was at 80.56E30, the stack was at 37.6E10, and the stash was at 181.2E10.

Monday June 11

The midnight shift began with Operations monitoring store 5484 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

At 2:22 AM, the MI power supply interlock loop (PSLOOP) tripped off due to a ground fault indication. Operations could not find a ground fault. They reset the loop.

At 7:02 AM, a Booster vacuum pump (BRFIP17) tripped off and caused BRF7 & 8 to trip. Everything reset.

Operations reported at 7:59 AM that the Recycler had stashed five times during the last twelve hours.

Operations terminated store 5484 at 8 AM.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinator will allow Pbar and MI experts to conduct studies.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/11/2007    email Fermilab